OpenStudy (anonymous):

Three consecutive integers have a sum of 42. Find the integers.

OpenStudy (texaschic101):

your consecutive integers are : x, x + 1, and x + 2 they equal 42... x + (x + 1) + (x + 2) = 42 --- combine like terms 3x + 3 = 42 -- subtract 3 from both sides 3x = 42 - 3 3x = 39 -- divide both sides by 3 x = 13 x + 1 = 13 + 1 = 14 x + 2 = 13 + 2 = 15 check.. 13 + 14 + 15 = 42 42 = 42 (correct) so your numbers are : 13,14,and 15